OpenStudy (anonymous):

GraPH THE QUADRATIC function label the vertex axis of symmetry and x intercepts also domain and range y=3x^2+21x+36 Graph included

OpenStudy (campbell_st):

take 3 out as a common factor y =3(x^2 + 7x + 12) factorise the quadratic 0 = 3(x+3)(x+4) so the intercepts: y intercept (x = 0) is 36 x intercepts (y = 0) x = -4, -3 axis of symmetry is x = -b/2a x = -21/6 x = -3.5 the axis if symmetry is x = -3.5 substitute it into the equation to find the vertex y = 3(-3.5)^2 + 21(-3.5) + 36 y = -0.75 vertex (-3.5, -0.75) domain is all real x range y >= -0.75